Solv for the indicated letter
b=3c, for c
The solution is c=_____

Found 2 solutions by DRichardson, stanbon:
Answer by DRichardson(63)

Hello There,
I just solved this equation for another student or maybe the same :)
c = b over three
Delilah r.

Answer by stanbon(75887)

Solve
for the indicated letter
b=3c, for c
Divide both sides by 3 to get:
b/3 = c
Interchange the sides.
c = b/3
The solution is c = b/3
============================
Cheers,
Stan H.